# SplAgger: Split Aggregation for Meta-Reinforcement Learning

### 08 Mar 2024 v2

@@ Abstract
A core ambition of reinforcement learning (RL) is the creation of agents capable of rapid learning in novel tasks. Meta-RL aims to achieve this by directly learning such agents. Black box methods do so by training off-the-shelf sequence models end-to-end. By contrast, task inference methods explicitly infer a posterior distribution over the unknown task, typically using distinct objectives and sequence models designed to enable task inference. Recent work has shown that task inference methods are not necessary for strong performance. However, it remains unclear whether task inference sequence models are beneficial even when task inference objectives are not. In this paper, we present strong evidence that task inference sequence models are still beneficial. In particular, we investigate sequence models with permutation invariant aggregation, which exploit the fact that, due to the Markov property, the task posterior does not depend on the order of data. We empirically confirm the advantage of permutation invariant sequence models without the use of task inference objectives. However, we also find, surprisingly, that there are multiple conditions under which permutation variance remains useful. Therefore, we propose SplAgger, which uses both permutation variant and invariant components to achieve the best of both worlds, outperforming all baselines on continuous control and memory environments.
@@ 3.2 Permutation Invariance
Figure 1: The hypernetwork from Beck et al. (2023b) is depicted in 0(a), the AMRL model from Beck et al. (2020) is depicted in 0(b), and SplAgger is depicted in 0(c). The angled line indicates a split connection that divides the neurons in half. The red arrow indicates a modified gradient computation in the backward pass. A hypernetwork is indicated by $h$. SplAgger makes use of the hypernetwork architecture combined with the AMRL sequence model. The hypernetwork architecture is necessary for performant end-to-end training. Critically, SplAgger also removes the gradient modification from AMRL which we show to be deleterious to performance.
@@ 4 SplAgger
Figure 2: A preview of later results. The permutation invariance of the max aggregator improves returns relative the RNN on the MC-LS environment (Beck et al., 2020), but decreases returns on the Planning Game (Ritter et al., 2021). Additionally, the ST gradient decreases the returns of the max aggregation. These results motivate SplAgger, which achieves the highest returns.
@@ 5 Experiments
On the four primary benchmark environments, we compare to four baselines. Since prior results demonstrate the need for hypernetworks when training end-to-end (Beck et al., 2023b), all baselines have been evaluated using hypernetworks, with design choices detailed in Appendix B. We additionally present negative results on a novel initialization method in Appendix E. The baselines evaluated primarily differ in their choice of aggregation function, $g$, and encoding of inputs, $e_{t}$. The baselines are described below.
@@ 6.2 Gradient Decay
Figure 8: An empirical analysis of gradients in SplAgger, AMRL, and RNNs. In 7(a) we see that the gradient of the output of the sequence model with respect to the initial input decreases over time. AMRL modifies the gradients to prevent this, but at the cost of exploding gradients with respect to the parameters, depicted in 7(b). We find that poor performance of the RNN is rather due to earlier inputs having a smaller gradient, when considering the final output of the sequence model. depicted in 7(c). Finally, models that are the least permutation variant do not necessarily perform better; the highest performing model, SplAgger, has an intermediate difference, but has components that are both permutation variant and permutation invariant. This is shown in 7(d).
@@ Appendix B Hyperparameter Tuning
(c) Walker
Figure 9: Tuning the KL-divergence weight on PEARL, in 8(a) and 8(b), and tuning the softmax temperatures in Figure 8(c). For PEARL, the chosen weight, $1^{-6}$, achieved the highest returns. Using a weight of 00 achieved a similar final returns, suggesting the stochastic latent was not particularly useful in our experiments. We chose $1^{-6}$ as it achieved the greatest average return throughout training, on both environments. For the softmax aggregators, 0.1 was the best initialization for the temperature.
We tune each baseline over five learning rates for the policy, [3e-3, 1e-3, 3e-4, 1e-4, 3e-5], for three seeds each (with a the confidence intervals (with a 68% confidence interval for all plots). Of the models evaluated, PEARL, and aggregators with the softmax aggregation, each require an additional hyperparameter. PEARL require a weight on objective penalizing the KL-divergence to the prior. We tune this weight on the T-LS and Walker environments, with a weight of $1^{-6}$ performing the best. In Figures 8(a) and 8(b), we display these results. Additionally, for softmax aggregators, we tuned the initial temperature of the softmax function. The temperature is learnable, however changes very slowly, and so is sensitive to initialization. This temperature was tuned between 1.0 and 0.1 on T-LS, with 0.1 performing the best. Note that we tuned the aggregators without using an RNN or SplAgger. This is depicted in Figure 8(c).
@@ Appendix C PEARL Design Choices and Additional Experiments
Figure 10: PEARL using the VariBad-stlye supervized reconstruction from Zintgraf et al. (2021) to train the sequence model (PEARL-vari). PEARL-vari, increases returns slightly on T-LS and Walker, but remains lower than SplAgger.
@@ Appendix D Analysis of PEARL Aggregation Failure
Figure 11: In 10(a), we see that the variance of PEARL decreases rapidly to zero at initialization. In 10(b), we see the same phenomenon, but when PEARL is fully trained.

### 01 Jun 2024 v3

@@ Abstract
A core ambition of reinforcement learning (RL) is the creation of agents capable of rapid learning in novel tasks. Meta-RL aims to achieve this by directly learning such agents. Black box methods do so by training off-the-shelf sequence models end-to-end. By contrast, task inference methods explicitly infer a posterior distribution over the unknown task, typically using distinct objectives and sequence models designed to enable task inference. Recent work has shown that task inference methods are not necessary for strong performance. However, it remains unclear whether task inference sequence models are beneficial even when task inference objectives are not. In this paper, we present evidence that task inference sequence models are indeed still beneficial. In particular, we investigate sequence models with permutation invariant aggregation, which exploit the fact that, due to the Markov property, the task posterior does not depend on the order of data. We empirically confirm the advantage of permutation invariant sequence models without the use of task inference objectives. However, we also find, surprisingly, that there are multiple conditions under which permutation variance remains useful. Therefore, we propose SplAgger, which uses both permutation variant and invariant components to achieve the best of both worlds, outperforming all baselines evaluated on continuous control and memory environments. Code is provided at https://github.com/jacooba/hyper.
@@ 3.2 Permutation Invariance
Figure 1: The hypernetwork from Beck et al. (2023b) is depicted in 1(a), the AMRL model from Beck et al. (2020) is depicted in 1(b), and SplAgger is depicted in 1(c). The angled line indicates a split connection that divides the neurons in half. The red arrow indicates a modified gradient computation in the backward pass. A hypernetwork is indicated by $h$. SplAgger makes use of the hypernetwork architecture combined with the AMRL sequence model. The hypernetwork architecture is necessary for performant end-to-end training. Critically, SplAgger also removes the gradient modification from AMRL which we show to be deleterious to performance.
@@ 4 SplAgger
Figure 2: A preview of later results. The permutation invariance of the max aggregator improves returns relative the RNN on the MC-LS environment (Beck et al., 2020), but decreases returns on the Planning Game (Ritter et al., 2021). Additionally, the ST gradient decreases the returns of the max aggregation. These results motivate SplAgger, which achieves the highest returns. (Results are reported with a 68% confidence interval, computed through bootstrapping with 1,000 iterations across three seeds, consistent with all plots presented.)
@@ 5 Experiments
On the four primary benchmark environments, we compare to four baselines. Hyperparamter tuning is detailed in Appendix B. Since prior results demonstrate the need for hypernetworks when training end-to-end (Beck et al., 2023b), all baselines have been evaluated using hypernetworks, with design choices detailed in Appendix B. We additionally present negative results on a novel initialization method in Appendix E. The baselines evaluated primarily differ in their choice of aggregation function, $g$, and encoding of inputs, $e_{t}$. The baselines are described below.
@@ 6.2 Gradient Decay
Figure 8: An empirical analysis of gradients in SplAgger, AMRL, and RNNs. In 8(a) we see that the gradient of the output of the sequence model with respect to the initial input decreases over time. AMRL modifies the gradients to prevent this, but at the cost of exploding gradients with respect to the parameters, depicted in 8(b). We find that poor performance of the RNN is rather due to earlier inputs having a smaller gradient, when considering the final output of the sequence model. depicted in 8(c). Finally, models that are the least permutation variant do not necessarily perform better; the highest performing model, SplAgger, has an intermediate difference, but has components that are both permutation variant and permutation invariant. This is shown in 8(d).
@@ Appendix B Hyperparameter Tuning
(c) T-LS
Figure 9: Tuning the KL-divergence weight on PEARL, in 9(a) and 9(b), and tuning the softmax temperatures in Figure 9(c). For PEARL, the chosen weight, $1^{-6}$, achieved the highest returns. Using a weight of 00 achieved a similar final returns, suggesting the stochastic latent was not particularly useful in our experiments. We chose $1^{-6}$ as it achieved the greatest average return throughout training, on both environments. For the softmax aggregators, 0.1 was the best initialization for the temperature.
We tune each baseline over five learning rates for the policy, [3e-3, 1e-3, 3e-4, 1e-4, 3e-5], for three seeds. (Results are reported with a 68% confidence interval, computed through bootstrapping with 1,000 iterations across three seeds, consistent with all plots presented.) Of the models evaluated, PEARL, and aggregators with the softmax aggregation, each require an additional hyperparameter. PEARL require a weight on objective penalizing the KL-divergence to the prior. We tune this weight on the T-LS and Walker environments, with a weight of $1^{-6}$ performing the best. In Figures 9(a) and 9(b), we display these results. Additionally, for softmax aggregators, we tuned the initial temperature of the softmax function. The temperature is learnable, however changes very slowly, and so is sensitive to initialization. This temperature was tuned between 1.0 and 0.1 on T-LS, with 0.1 performing the best. Note that we tuned the aggregators without using an RNN or SplAgger. This is depicted in Figure 9(c).
@@ Appendix C PEARL Design Choices and Additional Experiments
Figure 10: PEARL using the VariBad-stlye supervized reconstruction from Zintgraf et al. (2021) to train the sequence model (PEARL-vari). PEARL-vari, increases returns slightly on T-LS and Cheetah-Dir, but remains lower than SplAgger.
@@ Appendix D Analysis of PEARL Aggregation Failure
Figure 11: In 11(a), we see that the variance of PEARL decreases rapidly to zero at initialization. In 11(b), we see the same phenomenon, but when PEARL is fully trained.